OpenStudy (anonymous):

What is the simplified form of 9 times y to the seventh power over 15 times x to the fifth power divided by 5 x squared over 45 y cubed ?

OpenStudy (saifoo.khan):

\[\frac{9y^7}{15 x^5} \div \frac{x^2}{45y^3}\]

OpenStudy (saifoo.khan):

\[\frac{9y^7}{15 x^5} \times \frac{45y^3}{x^2}\]

OpenStudy (saifoo.khan):

\[\Large \frac{27y^{10}}{x^7}\]

OpenStudy (anonymous):

9y^7/15x^5 divided by 5x^/45y^3 like that

OpenStudy (saifoo.khan):

\[\frac{9y^7}{15 x^5} \div \frac{5x^2}{45y^3}\]

OpenStudy (saifoo.khan):

\[\frac{9y^7}{15 x^5} \times \frac{45y^3}{5x^2}\]

OpenStudy (anonymous):

the top one

OpenStudy (saifoo.khan):

\[\Large \frac{27y^{10}}{5x^7}\]